Solve the differential equation: (3- x?) y dy dx y +3/7–3
Answer #1

The given differential equation (3-2) y day de = y + 359-3 Y+ 3 ry a3 Joy du (3-2233 J-3=22 سے 723 Z+3+37 zde Now dy - azaz Yda Now (322) 3/2 ng Seo duaſ lost da 다 3 Cosodo 3-3 Sin 0 31 √ Cost 33 Cosmo S do 9 St Seu o do 1/3 tand Yv3 Coso Vir 3 3 3 3